# Negation: A Pink Elephant in the Large Language Models’ Room?

### v1

@@ Appendix B Full details of accuracy across evaluated models
Table 5: Evaluation of the negation sensitivity of each model across individual datasets. Our negation sensitivity metric represents the percentage of datasets in which the model assigned opposite entailment scores to pairs of hypotheses with and without negation. The best results (in bold) were achieved only by the biggest models supporting our RQ1.
Appendix C Prompting template
As a system prompt, we used the following template:
The user prompt follows a simple format:
Premise: <premise>
Hypothesis: <hypothesis>
The model decides by finishing a prepared response:
The answer is: <True / False>
Appendix D Statistical tests
In this section, we include the computation of the statistical tests upon which we base our claims in Section 3. Unless otherwise stated, for our computations, we are using alpha level (significance threshold) $\alpha=0.05$.
D.1 Model size and relative accuracy change
Using the results shown in Figure 4, we compute the correlation of the model size and the relative accuracy change per each evaluated dataset and overall. Using the Shapiro-Wilk test, we find that the majority of datasets do not have normal distribution over the different model sizes. Thus, we compute correlation using the Spearman correlation coefficient, as the Pearson correlation coefficient presumes a normal distribution of values. Values used during computation are shown in Table 6.
The results indicate a strong positive correlation between model size and the relative accuracy change on the FEVER dataset and a weak-to-moderate positive correlation on the SNLI dataset. For an averaged relative accuracy change for all languages per model, the results indicate a strong positive correlation with the Spearman correlation coefficient of $0.867$.
@@ D.1 Model size and relative accuracy change
D.1.1 Negation sensitivity
@@
D.2 Model family
As the model family is a categorical variable, instead of correlation, we compute and compare the variance of the relative accuracy change across the three model families (Qwen, Mistral, and Llama), using values shown in Figure 4 and the Kruskal-Wallis test as our analysis method. Our null hypothesis is that all model families have the same average of relative accuracy changes, and the alternative hypothesis is that the averages are not the same.
[…]
D.3 Language
We compute and compare the variance of the relative accuracy change across the four evaluated languages (CES, ENG, DEU, UKR), similarly to the model family variable. We aggregate values shown in Figure 4 per language and use the Kruskal-Wallis test as our analysis method. Our null hypothesis is that all languages have the same average of relative accuracy change, and the alternate hypothesis is that the averages are not the same.
[…]
As the $p$-value is not less than our alpha level, we cannot reject our null hypothesis. Thus, we conclude that there is no significant difference between the relative accuracy change across different languages.
D.4 Sentence complexity
As mentioned in Section 3, we posit that the larger models can more accurately capture nuanced language patterns, enabling them to correctly handle negations in more complex sentences that elude the smaller models. From each model family, we take the smallest and the largest models and compare the average sentence complexity of the hypotheses the model is able to correctly classify. However, we hypothesize that many hypotheses correctly classified by the smaller model are also correctly classified by the larger model. Therefore, for the larger model, we compute the average sentence complexity of the additional correctly classified hypotheses, i.e., the subset for the larger model constitutes of hypotheses the larger model correctly classified which are not present in the smaller model’s subset.
@@ D.4 Sentence complexity
Depth of the hypothesis dependency tree. Sentence complexity increases with the increase of the depth of the dependency tree.
Tables 8 and 9 show the the increase of average hypothesis depth between the smallest and the largest model per the language model family. Due to the relative simplicity of the hypotheses (short length as described in Table 2, average of one verb per hypothesis), the increase of depth is not substantial, being less than one level across all datasets for all language model families.
@@
Number of content words present only in the hypothesis and not the premise, divided by the total number of content words in the hypothesis. Sentence complexity increases with the increase of dissimilarity. We compute the similarity on the lemmatized forms of the words with the most common stop words per each language filtered out. High dissimilarity may be caused by the usage of synonyms, hypernyms, hyponyms, or other linguistic features that require higher understanding of syntactic patterns in the hypothesis.
Tables 10 and 11 show the increase in the average dissimilarity of the premise and its hypothesis between the smallest and the largest model per the language model family. Here we can see a marked increase in dissimilarity for hypotheses with negation, especially within the Qwen family models, showcasing the largest model’s increased ability in correctly reason and determine entailment on texts that use different formulations to express the same meaning.
@@
Qwen Family 0.038 -0.09 0.027 -0.016 0.065 -0.053 -0.084 -0.006 -0.080
Mistral Family -0.016 0.031 0.064 0.125 0.101 0.103 -0.012 0.151 -0.017
@@
Llama Family 2.52 % -0.12 % -4.31 % -5.77 % -15.30 % -4.73 % -16.01 % -1.20 % -6.83 %
Qwen Family -0.47 % 4.19 % 2.55 % 7.88 % -5.00 % 0.90 % 3.58 % 4.66 % 1.44 %
Mistral Family 0.89 % 4.49 % 0.42 % 6.94 % -1.63 % 1.02 % 6.26 % 2.49 % -1.92 %
Table 10: Increase of average sentence dissimilarity of premises and the evaluted hypotheses without negation that were correctly predicted only by the largest model of the LM family and the correctly predicted hypotheses by the smallest model of the LM family.

### v2

@@ 4 Results
Table 5: Evaluation of the negation sensitivity of each model across individual datasets, with the best results bolded.
Language plays a role in the models’ accuracy.
Figure 3 shows the accuracy change on different models with regard to the language of the dataset. As the language is a categorical variable, instead of correlation, we compute and compare the variance of the relative accuracy change across the four evaluated languages (CES, ENG, DEU, UKR), excluding the Czech manually translated dataset. We aggregate the values shown in Figure 4 per language and use the Friedman test as our analysis method. Our null hypothesis is that all languages have the same average of relative accuracy change, and the alternate hypothesis is that the averages are not the same.
ENG
DEU
UKR
CES 0.030 0.0273 0.0390
ENG 0.4961 0.4961
DEU 0.0195
ENG
DEU
UKR
0.039 0.6523 0.6523
0.2005 0.0977
0.2031
Table 6: $p$-values of Wilcoxon Signed-Test for language pairs on the SNLI (left) and the FEVER (right) datasets
[…]
As the $p$-value is less than our alpha level, we can reject our null hypothesis. Thus, we conclude that there is a significant difference in the relative accuracy change across different languages.
Delving further and applying the Wilcoxon Signed-Rank Test on the results of language pairs per specific dataset, we find that there is a statistically significant difference on the Czech SNLI dataset and all of the other languages, with the models performing worse on the Czech dataset. On the FEVER dataset, the models performed the worst on the Czech dataset as well, but the difference is statistically significant only in comparison to the English dataset.
Our results partially conform to our presuppositions in RQ2 that the models will perform worse in languages with higher word-order flexibility.
5 Conclusions and future work
This work presents a novel dataset enabling the evaluation of language models’ robustness in handling negations in four languages. We have extended existing FEVER Thorne et al. (2018) and SNLI Bowman et al. (2015) datasets to include paired hypotheses that are logically opposite to each other by negating the primary verb in the original hypothesis. We have evaluated the models’ accuracy separately on hypotheses with and without negation and computed the accuracy drop that occurred due to the addition of negation in the hypothesis. New datasets NoFEVER-ML and NoSNLI-ML are available in the anonymized dataset Anonymous (2025).
We find that model size plays a key role in handling negation correctly, countering the previous work of Truong et al. (2023). The results on differing language datasets only partially match the initial expectations, with models performing worse on the Czech configuration. This may be due to the relative simplicity of hypotheses, which do not lend to the inclusion of more complex negation structures, such as the negation token being farther away from the verb, as may be the case in German.
In regard to the differing accuracy from one dataset to another, we hypothesize that the accuracy is also heavily dependent on the specificity of information in the premise; the more common-sense SNLI dataset performed worse than the more fact-explicit FEVER dataset.
Our work provides new resources for addressing the deficiency of language models in handling negations. Our methodology and dataset can aid in finding the root cause of the negation problem by computing the accuracy on subsets of our new datasets in 4 languages, e.g., a subset of non-projective expressions of negation, a subset of different ways of expressing syntactic or semantic negation Kalouli et al. (2022), or to evaluate sub-clausal negation based on a dataset of Truong et al. (2022). We will use the lens of LLM-microscope Razzhigaev et al. (2025) to look at the tokens that play decisive rôle.
Limitations
The details of model preparation (training data, hyperparameters) limit the full understanding of the results we report, as we use pretrained models. Additionally, this paper does not delve into the effects of post-training and fine-tuning, and possible exposure to classes of negation examples as contradictory usage of negation. Also, the results might be biased by the semi-automated translations of datasets to their German, Czech, and Ukrainian versions, as seen in Table 4. However, we did a manual check to ensure the quality of a subset of the translations and did not encounter any errors related to the application of negation in context.
Appendix A Full details of accuracy across evaluated models
We include figures 6 and 6 as the full tables of accuracies of models for hypotheses with and without negation on the task of textual entailment. In both tables, we can see that almost all models achieve the highest accuracy on the English datasets, on both hypotheses, with and without negation.
@@ Appendix A Full details of accuracy across evaluated models
Appendix B Prompting template
As a system prompt, we used the following template:
The user prompt follows a simple format:
Premise: <premise>
Hypothesis: <hypothesis>
The model decides by finishing a prepared response:
The answer is: <True / False>
Appendix C Statistical tests
In tables 7 and 8 we provide the values used during statistical computations explained in Section 4.
@@ Appendix C Statistical tests
Appendix D Sentence complexity
As mentioned in Section 4, we posit that the larger models can more accurately capture nuanced language patterns, enabling them to correctly handle negations in more complex sentences that elude the smaller models. From each model family, we take the smallest and the largest models and compare the average sentence complexity of the hypotheses that the model is able to correctly classify. However, we hypothesize that many hypotheses correctly classified by the smaller model are also correctly classified by the larger model. Therefore, for the larger model, we compute the average sentence complexity of the additional correctly classified hypotheses, i.e., the subset for the larger model constitutes of hypotheses the larger model correctly classified which are not present in the smaller model’s subset.
@@ Appendix D Sentence complexity
Depth of the hypothesis dependency tree. Sentence complexity increases with the increase in the depth of the dependency tree.
Tables 9 and 10 show the increase of average hypothesis depth between the smallest and the largest model per the language model family. Due to the relative simplicity of the hypotheses (short length as described in Table 2, average of one verb per hypothesis), the increase of depth is not substantial, being less than one level across all datasets for all language model families.
@@
Number of content words present only in the hypothesis and not the premise, divided by the total number of content words in the hypothesis. Sentence complexity increases with the increase of dissimilarity. We compute the similarity on the lemmatized forms of the words, with the most common stop words per language filtered out. High dissimilarity may be caused by the usage of synonyms, hypernyms, hyponyms, or other linguistic features that require a higher understanding of syntactic patterns in the hypothesis.
Tables 11 and 12 show the increase in the average dissimilarity of the premise and its hypothesis between the smallest and the largest model per the language model family. Here we can see a marked increase in dissimilarity for hypotheses with negation, especially within the Qwen family models, showcasing the largest model’s increased ability to correctly reason and determine entailment on texts that use different formulations to express the same meaning.
@@
Qwen Family 0.038 --0.09 0.027 --0.016 0.065 --0.053 --0.084 --0.006 --0.080
Mistral Family --0.016 0.031 0.064 0.125 0.101 0.103 --0.012 0.151 --0.017
@@
Llama Family 2.52 % --0.12 % --4.31 % --5.77 % --15.30 % --4.73 % --16.01 % --1.20 % --6.83 %
Qwen Family --0.47 % 4.19 % 2.55 % 7.88 % --5.00 % 0.90 % 3.58 % 4.66 % 1.44 %
Mistral Family 0.89 % 4.49 % 0.42 % 6.94 % --1.63 % 1.02 % 6.26 % 2.49 % --1.92 %
Table 11: Increase of average sentence dissimilarity of premises and the evaluated hypotheses without negation that were correctly predicted only by the largest model of the LM family and the correctly predicted hypotheses by the smallest model of the LM family.